# Haonan Tang

### v3

@@ 1 Introduction
With the advent of deep learning, the TrackNet series [2019TrackNetv1, 2020TrackNetV2, 2023TrackNetV3p, 2023TrackNetV3up, 2024TrackNetV4] has emerged as the de facto benchmark in this domain. TrackNetV1 [2019TrackNetv1] pioneered the heatmap-based Encoder-Decoder paradigm but suffered from limited recall. TrackNetV2 [2020TrackNetV2] substantially improved throughput and accuracy via the Multiple-In-Multiple-Out (MIMO) architecture. More recently, TrackNetV3 [2023TrackNetV3up] demonstrated that stacking heavy computational modules (e.g., Inception-style blocks) could achieve higher accuracy, but at the cost of significantly increased latency (approx. $3\times$ slower than V2), rendering it less viable for resource-constrained real-time applications.
To balance efficiency and accuracy, the state-of-the-art TrackNetV4 [2024TrackNetV4] integrated motion priors based on absolute frame differences. While efficient, V4 and previous iterations still exhibit distinct structural limitations that hinder further performance breakthroughs:
@@ 2.1 Small Object Tracking in Sports
The TrackNet series [2019TrackNetv1, 2020TrackNetV2, 2023TrackNetV3p, 2023TrackNetV3up, 2024TrackNetV4] has established the dominant paradigm for high-speed ball tracking. TrackNetV1 [2019TrackNetv1] introduced the heatmap regression framework, treating tracking as a specific localization task. TrackNetV2 [2020TrackNetV2] advanced this by employing a Multiple-In-Multiple-Out (MIMO) architecture to process consecutive frames, significantly improving throughput. Subsequent variants, such as TrackNetV3 [2023TrackNetV3up], focused on maximizing accuracy by stacking heavyweight modules (e.g., Inception blocks), but at the cost of real-time performance.
Most recently, TrackNetV4 [2024TrackNetV4] integrated motion priors via absolute frame differences, successfully highlighting regions of movement. However, as discussed in Sec. 1, its reliance on absolute magnitudes inherently discards directional polarity, limiting the model’s ability to distinguish object trajectories from background noise. In contrast, our TrackNetV5 introduces a polarity-aware motion decoupling mechanism to recover these lost directional cues while maintaining high efficiency.
@@ 2.2 Motion Representation in Video Analysis
Explicit Motion Modeling. Methods like Optical Flow explicitly calculate pixel-level displacement vectors. Recent works, such as TOTNET [2025TOTNet], combine optical flow maps with 3D convolutions to extract robust spatio-temporal features. While accurate, optical flow computation incurs a prohibitive computational overhead, rendering it unsuitable for latency-sensitive applications.
Implicit Motion Modeling. Alternatively, architectures like I3D [2018I3Dv3] or Video Transformers [2021TimeSformer][2022VideoSwin] utilize 3D convolutions or self-attention to implicitly learn spatio-temporal dependencies. Although effective, these models typically require massive parameter counts and heavy computation (FLOPs).
@@ 2.3 Spatio-Temporal Refinement
Refinement strategies are widely used to correct initial predictions. In static keypoint detection, Cascaded Architectures (e.g., Convolutional Pose Machines [2016CPM]) refine heatmaps stage-by-stage. However, these spatial-only methods operate on single frames, ignoring the temporal consistency inherent in video data.
In the video domain, Temporal Attention mechanisms (e.g., TimeSformer [2021TimeSformer]) excel at capturing long-range dependencies. However, existing video Transformers are typically designed as heavyweight backbones rather than lightweight refinement heads. Furthermore, most refinement approaches employ a reconstruction paradigm, forcing the network to regenerate the entire heatmap from scratch.
@@ 3.1 Overall Architecture
Motion-Guided Feature Encoding. Given an input sequence of three consecutive frames $\mathcal{I}\in\mathbb{R}^{3\times 3\times H\times W}$, the processing initiates with the MDD module. As detailed in Section 3.1, MDD extracts directional polarity fields, which are immediately interleaved with the raw RGB frames to form a motion-augmented tensor $\mathcal{X}_{in}$. This composite input is fed into the V2 Backbone (detailed in Figure 2), which employs a U-Net[2015UNet]-like topology to capture multi-scale spatio-temporal contexts. The encoder progressively downsamples the feature space to a bottleneck representation, while the decoder reconstructs the spatial details via skip connections.
@@ 3.2.2 Learnable Attention Mapping
The raw pixel intensities in polarity maps $P^{+}$ and $P^{-}$ vary significantly with lighting conditions. To transform these signals into robust attention weights constrained to $[0,1]$, we employ a parameterized non-linear mapping function $f(\cdot)$. For a given polarity intensity $x$, the attention weight AA is computed as:
@@ 3.3.2 TSATTHead Architecture
Following the masking stage, we employ a patch embedding strategy [2021ViT] where non-overlapping patches are flattened and projected into an embedding space. These tokens are enriched with factorized spatio-temporal positional encodings to preserve sequence order. The sequence is then processed by a Transformer Encoder [2017Attention], where multi-head self-attention enables global interaction across spatial and temporal dimensions. Specifically, inspired by the factorized attention mechanism of TimeSformer [2021TimeSformer], we decompose the self-attention into separate spatial and temporal blocks to efficiently capture and reconstruct the trajectory coherence. Finally, the processed tokens are decoded via a PixelShuffle [2016ESPCN] operation to generate the high-resolution residual map $\Delta$, which is superposed onto the original clean draft to yield the final prediction.
@@ 4.1 Implementation Details
Datasets and Inputs. We evaluate our method on two datasets: the public TrackNetV2 Dataset ($1280\times 720$) and our internal Loveall Dataset ($1920\times 1080$). Following standard protocols [2020TrackNetV2], the data is split into training and validation sets with a 7:3 ratio. During training and inference, all input frames are resized to a resolution of $512\times 288$ to balance efficiency and spatial precision. The model takes a sequence of three consecutive RGB frames ($I_{t-1},I_{t},I_{t+1}$) as input to predict the trajectory heatmaps.
@@
Loss Function. Following TrackNetV2 [2020TrackNetV2], we utilize the Weighted Binary Cross Entropy (WBCE). This loss function addresses the extreme foreground-background imbalance by dynamically weighting the pixel-wise entropy based on prediction confidence. The formulation is defined as:
@@
Training Setup. The model is implemented in PyTorch and trained on a single NVIDIA RTX 4090 (24GB) GPU. We use the AdamW optimizer[2019AdamW] with a batch size of 2 and an initial learning rate of $1\times 10^{-4}$. The training spans 30 epochs, with a multi-step learning rate decay ($\gamma=0.1$) applied at epochs 20 and 25.

### v4

@@ 1 Introduction
With the advent of deep learning, the TrackNet series [6, 12, 4, 1, 9] has emerged as the de facto benchmark in this domain. TrackNetV1 [6] pioneered the heatmap-based Encoder-Decoder paradigm but suffered from limited recall. TrackNetV2 [12] substantially improved throughput and accuracy via the Multiple-In-Multiple-Out (MIMO) architecture. More recently, TrackNetV3 [1] demonstrated that stacking heavy computational modules (e.g., Inception-style blocks) could achieve higher accuracy, but at the cost of significantly increased latency (approx. $3\times$ slower than V2), rendering it less viable for resource-constrained real-time applications.
To balance efficiency and accuracy, the state-of-the-art TrackNetV4 [9] integrated motion priors based on absolute frame differences. While efficient, V4 and previous iterations still exhibit distinct structural limitations that hinder further performance breakthroughs:
@@ 2.1 Small Object Tracking in Sports
The TrackNet series [6, 12, 4, 1, 9] has established the dominant paradigm for high-speed ball tracking. TrackNetV1 [6] introduced the heatmap regression framework, treating tracking as a specific localization task. TrackNetV2 [12] advanced this by employing a Multiple-In-Multiple-Out (MIMO) architecture to process consecutive frames, significantly improving throughput. Subsequent variants, such as TrackNetV3 [1], focused on maximizing accuracy by stacking heavyweight modules (e.g., Inception blocks), but at the cost of real-time performance.
Most recently, TrackNetV4 [9] integrated motion priors via absolute frame differences, successfully highlighting regions of movement. However, as discussed in Sec. 1, its reliance on absolute magnitudes inherently discards directional polarity, limiting the model’s ability to distinguish object trajectories from background noise. In contrast, our TrackNetV5 introduces a polarity-aware motion decoupling mechanism to recover these lost directional cues while maintaining high efficiency.
@@ 2.2 Motion Representation in Video Analysis
Explicit Motion Modeling. Methods like Optical Flow explicitly calculate pixel-level displacement vectors. Recent works, such as TOTNET [15], combine optical flow maps with 3D convolutions to extract robust spatio-temporal features. While accurate, optical flow computation incurs a prohibitive computational overhead, rendering it unsuitable for latency-sensitive applications.
Implicit Motion Modeling. Alternatively, architectures like I3D [3] or Video Transformers [2][7] utilize 3D convolutions or self-attention to implicitly learn spatio-temporal dependencies. Although effective, these models typically require massive parameter counts and heavy computation (FLOPs).
@@ 2.3 Spatio-Temporal Refinement
Refinement strategies are widely used to correct initial predictions. In static keypoint detection, Cascaded Architectures (e.g., Convolutional Pose Machines [14]) refine heatmaps stage-by-stage. However, these spatial-only methods operate on single frames, ignoring the temporal consistency inherent in video data.
In the video domain, Temporal Attention mechanisms (e.g., TimeSformer [2]) excel at capturing long-range dependencies. However, existing video Transformers are typically designed as heavyweight backbones rather than lightweight refinement heads. Furthermore, most refinement approaches employ a reconstruction paradigm, forcing the network to regenerate the entire heatmap from scratch.
@@ 3.1 Overall Architecture
Motion-Guided Feature Encoding. Given an input sequence of three consecutive frames $\mathcal{I}\in\mathbb{R}^{3\times 3\times H\times W}$, the processing initiates with the MDD module. As detailed in Section 3.1, MDD extracts directional polarity fields, which are immediately interleaved with the raw RGB frames to form a motion-augmented tensor $\mathcal{X}_{in}$. This composite input is fed into the V2 Backbone (detailed in Figure 2), which employs a U-Net[10]-like topology to capture multi-scale spatio-temporal contexts. The encoder progressively downsamples the feature space to a bottleneck representation, while the decoder reconstructs the spatial details via skip connections.
@@ 3.2.2 Learnable Attention Mapping
The raw pixel intensities in polarity maps $P^{+}$ and $P^{-}$ vary significantly with lighting conditions. To transform these signals into robust attention weights constrained to $[0,1]$, we employ a parameterized non-linear mapping function $f(\cdot)$. For a given polarity intensity $x$, the attention weight $A$ is computed as:
@@ 3.3.2 TSATTHead Architecture
Following the masking stage, we employ a patch embedding strategy [5] where non-overlapping patches are flattened and projected into an embedding space. These tokens are enriched with factorized spatio-temporal positional encodings to preserve sequence order. The sequence is then processed by a Transformer Encoder [13], where multi-head self-attention enables global interaction across spatial and temporal dimensions. Specifically, inspired by the factorized attention mechanism of TimeSformer [2], we decompose the self-attention into separate spatial and temporal blocks to efficiently capture and reconstruct the trajectory coherence. Finally, the processed tokens are decoded via a PixelShuffle [11] operation to generate the high-resolution residual map $\Delta$, which is superposed onto the original clean draft to yield the final prediction.
@@ 4.1 Implementation Details
Datasets and Inputs. We evaluate our method on two datasets: the public TrackNetV2 Dataset ($1280\times 720$) and our internal Loveall Dataset ($1920\times 1080$). Following standard protocols [12], the data is split into training and validation sets with a 7:3 ratio. During training and inference, all input frames are resized to a resolution of $512\times 288$ to balance efficiency and spatial precision. The model takes a sequence of three consecutive RGB frames ($I_{t-1},I_{t},I_{t+1}$) as input to predict the trajectory heatmaps.
@@
Loss Function. Following TrackNetV2 [12], we utilize the Weighted Binary Cross Entropy (WBCE). This loss function addresses the extreme foreground-background imbalance by dynamically weighting the pixel-wise entropy based on prediction confidence. The formulation is defined as:
@@
Training Setup. The model is implemented in PyTorch and trained on a single NVIDIA RTX 4090 (24GB) GPU. We use the AdamW optimizer[8] with a batch size of 2 and an initial learning rate of $1\times 10^{-4}$. The training spans 30 epochs, with a multi-step learning rate decay ($\gamma=0.1$) applied at epochs 20 and 25.